OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution Awnser this?

OpenStudy (anonymous):

i think its d

OpenStudy (akashdeepdeb):

|x − 2| + 10 = 12 |x| can be +x or -x I have already told you WHY. :D So........ |x − 2| + 10 = 12 can be written in 2 ways. +(x-2) + 10 = 12 AND -(x-2) + 10 = 12 These will give you the max. and min. values for x. SOLVE THEM and you'll get the answer. Understood? :)

OpenStudy (anonymous):

just tellmethe answer instead.. @AkashdeepDeb

OpenStudy (akashdeepdeb):

I am not allowed to do that. And even if I was, I wouldn't, "For your own good." See the OpenStudy Code-of-Conduct @ http://www.openstudy.com/code-of-conduct I can help you instead. :)

OpenStudy (anonymous):

message me @AkashdeepDeb